2.1  A farming cooperative in the KwaMashu buys wheat seeds for its farmer members from seed merchants. A particular seed merchant claims that their wheat seeds have at least an 80% germination rate. Before the farming cooperative will buy from this seed merchant, they want to verify this claim. A random sample of 320 wheat seeds supplied by this seed merchant was tested, and it was found that only 230 seeds germinated. Is there sufficient statistical evidence at the 3% significance level to justify the purchase of wheat seeds from this seed merchant? Use the p-value approach to conduct a hypothesis test for a single proportion, and report the findings to the KwaMashu farming .            


1
Expert's answer
2021-10-06T12:16:21-0400

"\hat p=\frac{230}{320}=0.72."

"H_0:\pi=0.8."

"H_a:\pi>0.8."

Test statistic: "z=\frac{0.72-0.8}{\sqrt{\frac{0.8(1-0.8)}{320}}}=-3.58."

P-value: "p=P(Z>-3.58)=0.9998."

Since the p-value is greater than 0.03, fail to reject the null hypothesis.

There is no sufficient evidence that their wheat seeds have at least an 80% germination rate.